a wire is 224cm long is cut into three pieces the first is twice as long as the second the third is one-half as long as the second how long is the second piece?
Found 2 solutions by robertb, mananth:
Answer by robertb(5830)

f+s+t = 224,
f = 2s, and t =s/2. Direct substitution of the last two equations into the first gives
2s+s+s/2 = 224, or
7s/2 = 224, or
s = 64 cm, the length of the second piece.

Answer by mananth(16946)

a wire is 224cm long is cut into three pieces the first is twice as long as the second the third is one-half as long as the second how long is the second piece?
..
let second be x cm
first = 2x
third = x/2
..
x+2x+(x/2)=5x/2
7x/2=224
multiply by 2
(7x/2)*2 = 224*2
5x=448
/7
x=64
x=64 cm the second part
first part = 2x = 64*2=128 cm
third part = x/2 = 64/2=32 cm
...
m.ananth@hotmail.ca